OpenStudy (anonymous):

Write an equivalent fraction with the given denominator. 3=x/7. (I would like steps not just the answer)

OpenStudy (anonymous):

so do you need a fraction that has a denominator of 3 or 7?

OpenStudy (anonymous):

so the 3 needs to be written as a fraction with denominator 7?

OpenStudy (anonymous):

set up the proportion (two fractions equal to each other) as 3/1 = x/7

OpenStudy (anonymous):

now multiply the 1 times something to get a 7, so the something would be a 7...now do the same thing for the 3 in the numerator and 3 x 7 =21

OpenStudy (anonymous):

so 3 = 21/7 is this what you need?

OpenStudy (anonymous):

Yes thank you, its been helpful